### v1

# Low-resource Machine Translation: what for? who for? An observational study on a dedicated Tetun language translation service

Abstract
The impact of machine translation (MT) on low-resource languages remains poorly understood. In particular, observational studies of actual usage patterns are scarce. Such studies could provide valuable insights into user needs and behaviours, complementing survey-based methods. Here we present an observational analysis of real-world MT usage for Tetun, the lingua franca of Timor-Leste, using server logs from a widely-used MT service with over 70,0007000070,00070 , 000 monthly active users. Our analysis of 100,000100000100,000100 , 000 translation requests reveals patterns that challenge assumptions based on existing corpora. We find that users, many of them students on mobile devices, typically translate short texts into Tetun across diverse domains including science, healthcare, and daily life. This contrasts sharply with available Tetun corpora, which are dominated by news articles covering government and social issues. Our results suggest that MT systems for languages like Tetun should prioritise translating into the low-resource language, handling brief inputs effectively, and covering a wide range of domains relevant to educational contexts. More broadly, this study demonstrates how observational analysis can inform low-resource language technology development, by grounding research in practical community needs.
Low-resource Machine Translation: what for? who for?
An observational study on a dedicated Tetun language translation service
Raphael Merx1, Hanna Suominen2, Adérito José Guterres Correia3, Trevor Cohn1, Ekaterina Vylomova1 1School of Computing and Information Systems, The University of Melbourne 2The Australian National University 3Instituto Nacional de Linguística, Dili rmerx@student.unimelb.edu.au
1 Introduction
While machine translation (MT) is often considered a single area for natural language processing (NLP), it covers a large range of applications, end-user requirements, and language directions. MT can be leveraged for everyday communication, learning, or translation of official communication Dew et al. (2018); Vieira et al. (2021); Herrera-Espejel and Rach (2023); Paterson (2023). It can be used by students in the classroom Deng and Yu (2022), workers in an office environment Brynjolfsson et al. (2019), or by foreign travellers in public spaces Carvalho et al. (2023). Its reliability varies significantly depending on the size and domain coverage of available corpora Dabre et al. (2020); Khiu et al. (2024), and similarity between source and target languages Xu et al. (2020).
Figure 1: Real MT usage collides with assumptions made from available corpora
This disparity of potential applications and accuracy across domains or language pairs is particularly salient when translating to or from low-resource languages Haddow et al. (2022). In low-resource scenarios, the combination of high variability in MT accuracy across domains Haddow et al. (2022), and high variability of what native speakers value in NLP applications Lent et al. (2022), has prompted NLP researchers to better anchor their work in local community perception, including through surveys, focus groups, and product design Mager et al. (2023); Blaschke et al. (2024); Le Ferrand et al. (2022). These approaches, however, could be complemented by observational studies, that rely on end-user behaviours to infer needs or lack thereof for MT across domains or demographics.
In this paper on low-resource MT, we work with the Tetun language, which is the most spoken language in Timor-Leste General Directorate of Statistics Timor-Leste (2015). We rely on a sample of server logs from a volunteer-run MT service for the Tetun language, available in both website and app format. The service has over 70,0007000070,00070 , 000 monthly active users, and translates over one million documents per month. From a sample of 100,000100000100,000100 , 000 end-user translation text inputs, and their associated machine-translated output, we aim to respond to the following research questions:
  • Who are the main users of this service, what technology do they rely on, and what is the driver of their MT usage?
  • What characterises MT inputs, in terms of domain, length, direction of translation, and language pairs?
  • How do MT inputs compare to available corpora for the Tetun language?
2 Related Work
Tetun and the NLP community. While Tetun is a low-resource language de Jesus (2024), as of the 2015 census, it was the largest mother tongue in Timor-Leste with 361,027 speakers, and an additional 571,406 speakers of Tetun as a second language General Directorate of Statistics Timor-Leste (2015). It is the lingua franca in this country, spoken in most public spaces, and in the media Greksakova (2018). To our knowledge, the majority of Tetun NLP research to date was published by Gabriel de Jesus during his PhD candidature at the University of Porto. This included the creation of a language identification model, language tokenizer, and data pipeline for gathering Tetun monolingual data de Jesus and Nunes (2024a), the publication of that corpus de Jesus and Nunes (2024c), the study of retrieval strategies for Tetun text de Jesus (2024), and the study of different large language models (LLMs) relevance judgement capabilities in Tetun de Jesus and Nunes (2024b).
Community-driven approaches to NLP for low-resource languages. Recent years have witnessed a growing focus on developing NLP technologies for low-resource and endangered languages. However, researchers have identified that academic motivations can be at odds with the actual needs of language communities Le Ferrand et al. (2022); Liu et al. (2022), which made them advocate for a paradigm shift towards more community-centered approaches in NLP research Lent et al. (2022); Bird and Yibarbuk (2024). Studies emphasise the importance of close collaboration with language communities, understanding their specific needs, and involving them in the design and development process of language technologies Mager et al. (2023); Bird and Yibarbuk (2024). This approach acknowledges the heterogeneity among low-resource languages and their communities, recognising that methodologies effective for one may not be universally applicable Lent et al. (2022); Dolinska et al. (2024).
Linguistic analysis through large-scale data gathering. To our knowledge, this paper is the first large-scale analysis of real-world usage patterns in MT. It takes a similar approach to prior work leveraging large-scale secondary data in linguistics, including those predicting language learning difficulty Schepens et al. (2020), understanding what language characteristics lead to a language being confused for another Skirgård et al. (2017), and games that aim to incentivise the development of new language resources Fiumara et al. (2022).
3 Methodology
We work with data from a specialised Tetun MT service, which allows automated translation between Tetun Dili (tdt) and any of English (eng), Indonesian (ind), or Portuguese (por). The service also includes bilingual word dictionaries and a spell checker for the Tetun language, but we focus on the MT part. We get access to the following data sources:
  1. 1.
    A sample of 100,000100000100,000100 , 000 server logs from the machine translation service, all collected between March and August 2024. For each log, we get data on timestamp, translation input, MT output, source language code, requested target language code, and device OS.
  2. 2.
    Summary monthly active users’ data by country and device sourced from Google Analytics (for the website), Google Play Store (for the Android app), and Apple App Store (for the iOS app).
While data on the user base can be directly extracted from (2), domain analysis requires a choice of methodology that we describe in the following subsection.
3.1 Domain classification
Components of domain
Recognising that "domain" is often used interchangeably in NLP research to refer to topic (e.g. "the health domain") or source (e.g. "the news domain"), we follow the methodology in Saunders (2022), which defines domain as a combination of topic (one or more subject matters covered by the text), provenance (where the text comes from), and genre (language syntax and style).
3.1.1 Topic classification
LDA optimization
We start by identifying topics using Latent Dirichlet Allocation (LDA, Blei et al., 2003) on documents that are more than 5 words long, as very short documents do not offer enough context to identify their topic. We work with the Tetun side (either source Tetun text, or target machine-translated Tetun text), as this language is always present in any of the language pairs. The text is first lower cased and punctuation is removed. We run LDA enhanced with bigrams for improved context. After running hyperparameter search on number of topics, passes, and alpha, we find a maximum topic coherence of 0.065 using 15 topics, 10 passes, and an asymetric alpha. While this informs our choice of topic to categorise from, we qualitatively find that LDA is noisy for this task. For example, the health topic also includes words related to project management, such as “objetivu” or “servisu”. It also performs poorly on the categorisation of shorter documents, which are more likely to contain words that are not part of any identified topics.
Labels & test set preparation
Through a manual review of words for each of the LDA topic, we land on the following list of topics to classify from: Technology, Daily Life & Personal Experiences, Healthcare & Medicine, Law & Regulations, Business & Work & Employment, Government & Socio-economic Issues, Education, Science & Research, Religion & Spirituality. We map each of the 15 LDA topics to one of those topics, and we manually annotate a sample of 100 logs against this list of topics for evaluation, where each log can be classified against multiple topics.
Classification using pre-trained models
We evaluate other classification methods that rely on pre-trained models: zero-shot transformer-based classifiers and LLMs. Given the lack of pre-trained models that cover Tetun, we rely on the high-resource side of the input-output pair, which can be in English, Indonesian or Portuguese. For zero-shot classifiers, we use the HuggingFace zero shot classification pipeline [^1] https://huggingface.co/tasks/zero-shot-classification with the list of topics passed as candidate labels, in a multi-label mode. We evaluate bart-large-mnli [^2] https://huggingface.co/facebook/bart-large-mnli and deberta-v3-large-zeroshot-v2.0333https://huggingface.co/MoritzLaurer/deberta-v3-large-zeroshot-v2.0 Laurer and Lochner (2024). For LLMs, we evaluate LLama 3.1 8b Dubey et al. (2024), with a prompt that lists the domains to classify from, and that includes examples for in-context learning, shown in Appendix A.
Classification evaluation
We compare micro and weighted average scores on our test set for the different classification approaches in Table 1. We find that Llama 3.1 8B has the highest F1 score on both metrics for this classification task, followed by (in order) deberta-v3-large-zeroshot-v2.0, bart-large-mnli, and last LDA. We therefore pick Llama 3.1 8B for domain classification, with results reported in Section 4.3.
Approach Micro avg F1 Weighted avg F1
LDA $0.210.21$ $0.190.19$
bart-mnli4442 $0.400.40$ $0.400.40$
deberta-nli5553 $0.660.66$ $0.680.68$
Llama 3.1 8B $0.760.76$ $0.760.76$
Table 1: Comparison of micro and weighted average F1 score for different domain classification approaches. LDA is run on the Tetun side, while other approaches are run on the high-resource side (English/Portuguese/Indonesian).
3.1.2 Provenance classification
After a manual review of the text samples used in topic classification, we propose the following categories for provenance: Education & research material, News article & press release, Organisational & formal documents, Conversation & correspondence, Literary fiction, Religious text, Other. We use Llama 3.1 8B here as well, with a prompt that lists the sources to classify from, and that includes examples for in-context learning. This approach results in a weighted F1 score of 0.73 on our test set, which we deem satisfactory for classifying provenance, given the natural ambiguity of this task, especially for shorter documents. We report our results in Section 4.3.2.
3.1.3 Genre
After recommendation from the Timor-Leste Instituto Nacional de Linguística (INL, National Institute of Linguistics), under genre we evaluate what fraction of Tetun MT inputs follow official spelling, as defined by the country’s Matadalan Ortográfiku ba Tetun Nasionál (Orthographic Guide for National Tetun, 2002).
For calculating the percentage of words that follow INL spelling in each document, we apply the following preprocessing: (1) tokenize Tetun document using an open source Tetun tokenizer de Jesus and Nunes (2024a), (2) ignore words that start with a capital letter as they are often proper nouns and (3) ignore tokens that start with a number or punctuation character. We report our results in Section 4.3.3. We further ignore inputs where less than 20% of words are in the official spelling guide, as we empirically find that these documents are almost never in Tetun.
3.1.4 Comparison with available monolingual corpora
For each of the identified domain components (topic, provenance, genre), we provide a comparison between MT inputs and available Tetun corpora. For the latter, we select the Tetun "clean" split of the MADLAD-400 dataset Kudugunta et al. (2023), as it is the largest Tetun corpus available to date with 40,3674036740,36740 , 367 documents. For classifying topic and provenance, given that we work with models that do not support Tetun, we first translate the MADLAD Tetun corpus to English using the MADLAD-3B model. We release our translation of this corpus for future researchers to use, which includes 884k Tetun sentences and their machine-translated text.
For all domain analysis, we weigh inputs by their number of words. This is to ensure that very short MT inputs do not get as much weight as longer, more informative inputs, or that users that tend to translate several paragraphs at a time are not under-represented compared to users who translate shorter inputs.
4 Results
4.1 User base
Device. Most users access the service through their mobile device, either through the Android application (71,5007150071,50071 , 500 monthly active users), the iOS application (2,50025002,5002 , 500 monthly active users), or the website in a mobile browser (13,0001300013,00013 , 000 monthly active users). Desktop users of the website represented 8,30083008,3008 , 300 monthly active users, with an unknown number of these users being duplicate of mobile users. The distribution of users is illustrated in Figure 2. We discuss the implications of this high representation of mobile devices, particularly Android devices, on NLP research, in Section 5.
MobileDesktop00 k20202020 k40404040 k60606060 k80808080 kGoogle Play StoreApple App StoreWebsite
Figure 2: Monthly active users by device and service. Note that users who use both the website and mobile app would be double counted.
Country. Most users are in Timor-Leste: 67,000 of 71,500 (93.4%) Android monthly active users have their device registered in Timor-Leste, and 17,000 of 22,000 (77%) website monthly active users are in Timor-Leste. These numbers can be noisy: for example, Timorese users who live abroad might have a mobile device registered outside of Timor-Leste, and foreigners in Timor-Leste using the website cannot be differentiated from local users. We think, however, that the results support the hypothesis that the large majority of users are Timorese with English as a second language (ESL), as opposed to foreigners learning Tetun.
Demographics. Server usage patterns support the hypothesis that most users are students who use the service for educational purposes. A large proportion of requests happen in the early morning or in the evening (see Figure 3), at times when office workers are less likely to use the service, and when students are more likely to do their homework. Authors of the MT service also communicated to us that service usage goes down during school holidays, and back up on the evening before school resumes, another pattern that supports this hypothesis.
004444888812121212161616162020202000 k0.50.50.50.5 k1111 k1.51.51.51.5 k2222 kHour of day
Figure 3: Histogram showing number of requests by time of day, in Asia/Dili (UTC+9). We observe a spike in the evenings, when students are more likely to prepare their homework.
4.2 Translation direction and text length
Translation direction
Using the source and target language codes present in each of the server logs, we extract summary information on the distribution by source and target language. We find that Tetun is more likely to be the target language (68% of requests), with English to Tetun translation most in demand (48% of requests), followed by Tetun to English (25% of requests), Portuguese to Tetun (11%), Indonesian to Tetun (9%), Tetun to Portuguese (4.6%), and Tetun to Indonesian (2.6%) – see Table 2.
Direction Requests (% of total)
English to Tetun 48%
Portuguese to Tetun 11%
Indonesian to Tetun 9%
Total Tetun as Target 68%
Tetun to English 25%
Tetun to Portuguese 4.6%
Tetun to Indonesian 2.6%
Total Tetun as Source 32%
Table 2: Distribution of translation requests by language pair
Length of translated text
Leveraging an open source Tetun tokenizer de Jesus and Nunes (2024a) for Tetun, and the NLTK library for English/Indonesian/Portuguese, we count the number of words in the input texts. We find that shorter text inputs are much more common, with a sharp decline in frequency as the text length increases. Overall, inputted text has a median of 8 words, with a median of 12 words when Tetun is the target of translation, and 5 words when Tetun is the source. A histogram of word count frequency is shown in Figure 4.
025507500 k10101010 k20202020 k30303030 kLength of input text (number of words)
Figure 4: Histogram showing the frequency distribution of input text lengths measured by the number of words.
4.3 Domain coverage
4.3.1 Topic
Using the input text and its associated MT-translated result from the server logs, we estimate what topics are most or least in-demand for text translation. Our results for topic coverage are shown in Table 3. We include a comparison with topic coverage for the largest open Tetun dataset currently available, MADLAD-400 Kudugunta et al. (2023).
We find overall that "Science & research" is the topic most in demand, present in 34.2% of MT inputs when weighted by word count, followed by "Healthcare & medicine", "Education", and "Business & work & employment", all present in 20 to 25% of inputs. Overall, "Religion & spirituality" is the least frequent topic in MT inputs, which comes in contrast with the high representation of the religious domain in low-resource corpora Haddow et al. (2022).
We observe a large discrepancy between user-translated text and available corpora, where the main topics covered by available corpora (Government & socio-economic issues, 83.1%; Law & regulations, 28.2%) do not match with the main topics in demand by MT users.
Topic MT text MADLAD
Science & research 34.17% 3.173.173.173.17%
Healthcare & medicine 23.90% 11.9011.9011.9011.90%
Education 22.82% 11.2811.2811.2811.28%
Business & work & empl.. 21.19% 6.186.186.186.18%
Daily life & personal exp. 17.1517.1517.1517.15% 8.668.668.668.66%
Government & socio. issues 16.1716.1716.1716.17% 83.09%
Law & regulations 8.908.908.908.90% 28.21%
Technology 8.068.068.068.06% 3.903.903.903.90%
Religion & spirituality 6.806.806.806.80% 14.5114.5114.5114.51%
Table 3: Topic distribution, where each document can cover several topics. Highlighted values indicate percentages above 20%.
0020202020404040406060606080808080100100100100TechnologyDaily Life & Personal ExperiencesHealthcare & MedicineLaw & RegulationsBusiness & Work & EmploymentGovernment & Socio-economic IssuesEducationScience & ResearchReligion & Spirituality3.93.93.93.98.668.668.668.6611.911.911.911.928.2128.2128.2128.216.186.186.186.1883.0983.0983.0983.0911.2811.2811.2811.283.173.173.173.1714.5114.5114.5114.518.068.068.068.0617.1517.1517.1517.1523.923.923.923.98.98.98.98.921.1921.1921.1921.1916.1716.1716.1716.1722.8222.8222.8222.8234.1734.1734.1734.176.86.86.86.8Percentage (%)MT textMADLAD
Figure 5: Comparison of topic coverage for text translated on the MT service with the largest openly available Tetun corpus (MADLAD-400 "clean" split)
4.3.2 Provenance
See Table 4. We observe a large dominance of "Education & research material", which alone represent more than half of MT inputs (55.9%). "Organisational & formal documents" come as a distant second at 18%, followed by "Conversation & correspondence" at 8.7%. This again stands in contrast with monolingual corpora, where over 61% of documents come from "News articles & press releases", while only 7.5% come from "Education & research material".
Overall, our findings on topic and provenance are consistent with the hypothesis that most users of this MT service are students (who are more likely than the general population to translate educational material related to science, education or healthcare) and with de Jesus and Nunes (2024c) who find that most Tetun text on the internet comes from the news domain, which is likely to overlap with the "Government & socio-economic issues" topic.
Provenance MT text MADLAD
Education & research material 55.90% 7.517.517.517.51%
Organisational & formal docs 18.0018.0018.0018.00% 17.2817.2817.2817.28%
News article & press release 10.2010.2010.2010.20% 61.16%
Conversation & correspond. 8.728.728.728.72% 4.754.754.754.75%
Literary fiction 4.264.264.264.26% 0.890.890.890.89%
Religious text 2.522.522.522.52% 8.158.158.158.15%
Other 0.400.400.400.40% 0.260.260.260.26%
Table 4: Provenance distribution in MT and MADLAD (highlighted values indicate percentages above 20%)
4.3.3 Genre
Overall, we observe that official (INL) spelling is only loosely followed by MT inputs, with a median of 71% of words spelled correctly, compared to a median of 82% in MADLAD. This is coherent with previous studies that note the high proportion of Tetun texts that do not follow the official spelling, including in the media Greksakova (2018). We also observe that variance in the rate of correct spelling is higher in MT inputs (standard deviation of 13%) than in MADLAD (standard deviation of 8%) – see Figure 6.
25%50%75%100%MT Tetun inputs25%50%75%MADLAD
Figure 6: Histograms of MT Tetun inputs and MADLAD Tetun texts by fraction of words in the official INL spelling guide.
4.4 Further comparison with monolingual corpora
Surprisal
To illustrate the lack of domain overlap between MT inputs and available corpora identified in Section 4.3, we leverage a pre-trained Tetun language model to compare surprisal on MT Tetun inputs with surprisal on monolingual corpora. We use the Goldfish Tetun model Chang et al. (2024), which is trained mostly (> 98%) on MADLAD Tetun data. We get an average surprisal of 6.1 on MT Tetun inputs, compared to an average of 4.3 for Labadain (where we exclude sentences that appear in MADLAD). This higher surprisal suggests that the MT inputs may contain vocabulary, structures, or contexts that are underrepresented in scraped monolingual corpora, potentially leading to lower quality translations or increased translation errors.
Comparing average surprisal per domain, we find that documents in the "Government & Socio-economic issues" and "Religion & spirituality" domains have a lower surprisal (5.74 and 5.86 respectively) than documents from the "Science & research" and Technology (6.42 and 6.69 respectively). This result can be explained by the domain mismatch between MT inputs and monolingual corpora uncovered in Section 4.3.
Semantic embeddings
Working with a sample of 1,000 translation logs and 1,000 MADLAD Tetun documents, we aim to get a semantic representation of documents for comparison. Given the lack of Tetun semantic embedding model, we work with the high-resource side (in English, Indonesian or Portuguese) for translation logs, and we translate the MADLAD Tetun data to English using MADLAD-3B. We use LASER encoders Team et al. (2022), which support English, Indonesian and Portuguese, to get a shared representation of both translation logs and translated MADLAD data. After embedding, we visualise our result using t-SNE Van der Maaten and Hinton (2008). We find overall a large degree of separation between translation logs and MADLAD monolingual Tetun corpora, shown in Figure 7.
Figure 7: t-SNE visual comparison of MT text vs MADLAD embeddings
5 Recommendations
Future NLP work. Our results seem to indicate that future work on Tetun MT should focus on: (1) broad domain coverage that would include texts about daily life, science and health, (2) accuracy when translating into Tetun rather than from Tetun, and (3) accuracy when translating very short documents (including individual words). This could start with an evaluation framework that finds a balance between these criteria. In particular, the sharp imbalance between the domain of available Tetun corpora and MT inputs leads us to recommend the systematic evaluation of MT accuracy in a wide variety of domains, and advise against measuring MT accuracy on domains that are over-represented in available corpora, as this would risk giving a poor representation of end-user experience.
Future Human-Computer Interaction (HCI) work. We recommend a focus on meeting the needs of students who use MT on their mobile device. Given that students seem to represent a large portion of MT users, and that learning outcomes are an important social good, future study of MT perception and improvement for students in low-resource contexts would be beneficial, both to the research community and end users. Similarly, given the high cost and low reliability of internet connections in Timor Foundation (2023), we recommend a heightened focus on MT inference on mobile devices, for example through further research into efficient MT, or through the creation of an open source framework for on-device MT that targets mobile devices [^6] Similar to the Bergamot / Browser MT project.
6 Limitations
Design decisions made by the volunteers who built this MT service might have impacted usage numbers. For example, apart from MT, the service includes a multilingual word dictionary (with entry definitions and example sentences), and a Tetun spell checker, which might skew the number of users upward, with potentially some users using the app or website without using the MT functionality. Similarly, the authors chose to only allow translation between Tetun and three higher-resource languages (English, Portuguese, Indonesian), which raises the question of how adding support for more languages would impact usage and balance between languages.
This paper focuses on Tetun, which might limit the applicability of our findings to other languages, given the variability of use cases, users needs and data availability within low-resource languages Liu et al. (2022). While some of our findings will be applicable to other low-resource languages, in particular those that are also lingua franca, they are less likely to be applicable to much lower-resourced languages, in particular endangered ones.
Lastly, we identify some potential shortcomings of our domain classification methodology: (1) some classes are overlapping (e.g. text about biology could fall under "Health & medicine" or "Science & technology"), which makes them harder to distinguish, (2) it does not identify the source of each text, e.g. a document about healthcare could be from a news article or from a research article, and (3) it works only on documents that longer than 5 words, but most translated documents are very short, thus not covered by our classification methodology.
7 Conclusion
We present an analysis of MT usage for Tetun, a low-resource language, based on server logs from a widely-used MT service. Our study reveals key insights into user demographics, device preferences, translation directions, and domain coverage. We find that a large share of users are likely students in Timor-Leste using mobile devices, with a preference for translating into Tetun rather than from it. Our analysis also uncovers a significant mismatch between the domains of user-translated text and existing Tetun corpora.
Our recommendations include a stronger emphasis on domains that are not well covered by existing corpora (in particular domains that are important to education), accuracy for translation into Tetun, performance on very short texts, and the development of mobile-optimised MT solutions. For the HCI community, we suggest focusing on MT perception and improvement for students in low-resource contexts.
This study provides a first observational analysis of real-world MT usage for a low-resource language. It contributes to a more nuanced understanding of the challenges and opportunities in this field. We hope these insights will guide future research and development efforts, ultimately leading to more effective and accessible MT solutions for speakers of low-resource languages.
Appendix A Prompt for domain classification
TEXT_PROMPT = """<texttoclassify>
{text}
</texttoclassify>"""
EXAMPLE_USER_1 = TEXT_PROMPT.format(text="Pleasemeetmeattheparkat3pm.")
EXAMPLE_RESPONSE_1 = """<rationale>
Thistextisaboutgoingtotheparkwithafriend,asituationthatispartofdailylifeexperiences.
</rationale>
<domain>
dailylife&personalexperiences
</domain>"""
EXAMPLE_USER_2 = TEXT_PROMPT.format(text="Mindfulness-basedinterventions(MBIs)areincreasinglybeingintegratedintooncologicaltreatmenttomitigatepsychologicaldistressandpromoteemotionalandphysicalwell-being.")
EXAMPLE_RESPONSE_2 = """<rationale>
Thistextisabouttheuseofmindfulness-basedinterventionsinoncologicaltreatment,ahealthtopic.Itseemstobefromascientificorresearchcontext.
</rationale>
<domain>
healthcare&medicine;science&research
</domain>"""
EXAMPLE_USER_3 = TEXT_PROMPT.format(text="Thegovernmenthasannouncednewpoliciestoboostemploymentopportunitiesinruralareas.")
EXAMPLE_RESPONSE_3 = """<rationale>
Thistextdiscussesgovernmentpoliciesaimedatimprovingemployment,whichrelatestobothgovernanceandsocio-economicissues,aswellasbusinessandemployment.
</rationale>
<domain>
government&socio-economicissues;business&work&employment
</domain>"""
EXAMPLE_USER_4 = TEXT_PROMPT.format(text="Neracaadalahsuatulaporankeuanganjadididalamnyapastiterdapattigabagianterpentingyaituasetatauharta,liabilitasatauutang,danekuitas.")
EXAMPLE_RESPONSE_4 = """<rationale>
Thistextdescribesafinancialreport,specificallythebalancesheet,whichisanessentialcomponentofaccountingandfinance.Itbelongstothebusinessandfinancedomain.
</rationale>
<domain>
business&work&employment
</domain>"""
messages = [
{"role": "system", "content": f"Youarealinguistassistant,helpingclassifyuser-inputtedtextintoalistofdomains:<listofdomains>{’;’.join(TOPICS)}</listofdomains>.Youwillbegiventexttoclassify,andaskedthedomainsthatthistextbelongsto.Onlyeverusethedomainslistedin<listofdomains>"},
{"role": "user", "content": EXAMPLE_USER_1},
{"role": "system", "content": EXAMPLE_RESPONSE_1},
{"role": "user", "content": EXAMPLE_USER_2},
{"role": "system", "content": EXAMPLE_RESPONSE_2},
{"role": "user", "content": EXAMPLE_USER_3},
{"role": "system", "content": EXAMPLE_RESPONSE_3},
{"role": "user", "content": EXAMPLE_USER_4},
{"role": "system", "content": EXAMPLE_RESPONSE_4},